OpenStudy (anonymous):

Find S25 for -7+ -2 + 3 + 8 +… I got -2275 following but I don't think that's right. Any advice would be awesome! Thanks.

OpenStudy (anonymous):

How'd you get that?

OpenStudy (anonymous):

Wouldn't it just be 113?

OpenStudy (anonymous):

so we know that a1 = -7 tell me what is d ? d = a2 - a1

OpenStudy (anonymous):

when you find d you can find a25: a25 = a1 + (25-1)*d = a1 + 24d

OpenStudy (anonymous):

after you find a25 you can find S25 : S25 = 25 * (a1 + a25) / 2

OpenStudy (anonymous):

@Coolsector : Your answer is right !:) :D @maddie.jones47563 : Flow the Coolsector s answer !

OpenStudy (anonymous):

Okay. Thanks you guys :) @pathosdebater @E.ali @Coolsector